OpenStudy (anonymous):

Which fraction falls between three-tenths and three-fifths? five-tenths eight-tenths four-fifths two-tenths

OpenStudy (opcode):

three-tenths and three-fifths: 3/10 and 3/5. Get a common denominator.

OpenStudy (anonymous):

3/5=6/10

OpenStudy (opcode):

\[\dfrac{3}{10}\] \[\dfrac{3}{5}\times\dfrac{2}{2}\Rightarrow\dfrac{6}{10}\]

OpenStudy (opcode):

So: Five-tenths \[\Rightarrow \dfrac{5}{10}\]